OpenStudy (anonymous):

suppose that a 50lb box sits on a ramp that makes a 20 degree angle with the horizontal a. find the component of the 50 lb force that is parallel to the ramp... I have no idea what this question is asking. b. which component, horizontal or vertical moves the sled along the gorund

OpenStudy (accidentalaichan):

Okay, let's give this a shot The 50 would be the hypotenuse of a right triangle with the parallel to the ramp vector (r) an opposite leg from the 20º angle so by soh cah toa you'd use sin 20 = r/50 r = 50 sin 20 = 17.10 lb The perpendicular to ramp vector p is the adjacent side so cos 20 = p/50 p = 50 cos 20 = 46.98 lb

OpenStudy (accidentalaichan):

did it help?

OpenStudy (anonymous):

yes thank you